Percent of Sales Method At the end of the current year, Accounts Receivable has a balance of $835,000; Allowance for Doubtful Accounts has a debit balance of $7,500; and sales for the year total $3,760,000. Bad debt expense is estimated at 1/4 of 1% of sales.

1. Determine the amount of the adjusting entry for uncollectible accounts. $

2. Determine the adjusted balances of Accounts Receivable, Allowance for Doubtful Accounts, and Bad Debt Expense.

Accounts Receivable $

Allowance for Doubtful Accounts $

Bad Debt Expense $

3. Determine the net realizable value of accounts receivable. $
Answer #1

1) Adjusting entry = 3760000*1%*1/4 = 9400

2) Bad debt expense = 9400

Account receivable = 835000

Allowance for doubtful accounts = 7500-9400 = 1900

3) Net realizable value = 835000-1900 = 833100